OpenStudy (anonymous):

please help medaling and fanning Which is not a solution to the equation ? A. (–4, –7) B. (3, 14) C. (2, –1) D. (1, 8)

OpenStudy (a1234):

Equation?

OpenStudy (anonymous):

oh sorry -3x + y = 5

OpenStudy (a1234):

Use the x- and y-values in the choices and put them into the equation. I'll show you one. -3(-4) + (-7) = 5 12 - 7 = 5 5 = 5 So, (-4, -7) is a solution to the equation.